Given

α=i+2j+3k,

β=2i+j−k find

α×β


1
Expert's answer
2020-06-15T19:40:18-0400

Solution.

"\u03b1=i+2j+3k;"

"\u03b2=2i+j\u2212k";

"α×β=\begin{vmatrix}
 i & j&k \\
 1& 2&3\\
 2&1&-1
 \end{vmatrix}=i(2\sdot(-1)-1\sdot3)-j(1\sdot(-1)-2\sdot3)+"

"+k(1\sdot1-2\sdot2)=i(-5)-j(-7)+k(-3)=-5i+7j-3k;"

Answer: "\u03b1\u00d7\u03b2=-5i+7j-3k."